Calculate the nominal and real returns as well as the nominal and real risk premiums for the following corporate bond investment: Purchased for $940 one year ago, 4% coupon rate, sold for $994. The inflation rate was 5.0% and T-bills returned 6%
Answer #1

Sol)

Considering face value of $1000

Nominal rate of return = (coupon interest + Bond appreciation)/(beginning price) *100

= (40 + 54)/(940) *100

= 94/940 *100

= 10%

Real rate of return = ((1+ nominal rate)/(1+inflation))-1

= (1.10/1.05) -1

=4.76%